1. Find the slope-intercept equation of the line passing through (-2, -6) which is also
a. parallel to 4x-16y=2
b. perpendicular to 4x-16y=2

Answer by mananth(16949)

1. 4x-16y=2
16y=4x-2
/16
y=(1/4)x-(1/8)
slope of this line is (1/4)
The parallel line will have same slope = 1/4
point = (-2,-6)
Find b , plug point and slope in the equation y = mx+b
-6=(1/4)(-2)+b
-6=(-1/2)+b
-6+1/2=b
b=-5 1/2
b= -11/2
Equation ; y= (1/4)x-11/2
---
4 x -16 y = 2
Find the slope of this line

-16 y = -4 x + 2
Divide by -16
y = 0.250 x + -0.13
Compare this equation with y=mx+b
slope m = 1/4

The slope of a line perpendicular to the above line will be the negative reciprocal -4
m1*m2=-1
The slope of the required line will be -4
m= -4 ,point ( -2 , -6 )
Find b by plugging the values of m & the point in
y=mx+b
-6 = 8 + b
b= -14
m= -4
The required equation is y =-4x-14
m.ananth@hotmail.ca